# Surface Reconstruction Using Rotation Systems

### v1

@@ Abstract.
Inspired by the seminal result that a graph and an associated rotation system uniquely determine the topology of a closed manifold, we propose a combinatorial method for reconstruction of surfaces from points. Our method constructs a spanning tree and a rotation system. Since the tree is trivially a planar graph, its rotation system determines a genus zero surface with a single face which we proceed to incrementally refine by inserting edges to split faces and thus merging them. In order to raise the genus, special handles are added by inserting edges between different faces and thus merging them. We apply our method to a wide range of input point clouds in order to investigate its effectiveness, and we compare our method to several other surface reconstruction methods. We find that our method offers better control over outlier classification, i.e. which points to include in the reconstructed surface, and also more control over the topology of the reconstructed surface.
@@
copyright: none
@@ 1. Introduction
The ill-posed nature of the problem is due to the fact that there is generally no unique solution. Even if we restrict ourselves to only consider methods which connect all points to form a 2-manifold orientable triangle mesh, the number of possible solutions is immense for realistic input sizes. While it is trivial to connect a point to its $k$ nearest neighbors to form a graph, it is far from trivial to form a triangle mesh from this collection of edges, precisely due to the aforementioned combinatorial complexity. However, if we require the output triangle mesh to have disk topology, graph theory provides a solution, that to the best of our knowledge has not been considered before. The restriction to disk topology would be problematic, but we show how it is possible to relax this restriction in order to obtain meshes with boundary curves and of arbitrary genus.
@@
The method to which we allude is based on the simple observation that a tree (in the graph theoretical sense) is always a planar graph, and the planar embedding of a tree is given by a rotation system which defines the clockwise ordering of edges incident on each vertex. Thus, if we have a spanning tree connecting all of our points and an associated rotation system, we effectively have a polygonization with a single polygon whose edges are simply the edges of the tree (with each edge of the tree appearing twice). To obtain a triangle mesh, we recursively insert edges that split polygons into smaller polygons. The process stops when all polygons - except the boundary polygon - are triangles. A simple 2D example is shown in Figure 4 for a synthetic point cloud.
@@
For moderate input sizes and a 2D point cloud, the algorithm outlined above is simple to implement because the points are already embedded in the plane, and the rotation system is given implicitly. On the other hand, when designing an efficient and effective 3D version of the algorithm, we face several challenges. First of all, the rotation system, i.e. the ordering of outgoing edges for each vertex, must be computed consistently for all vertices. Moreover, when connecting points, several geometric considerations are important, and we want to allow connections that violate planarity in order to obtain meshes of arbitrary genus. Finally, efficiency is a concern: when an edge is added, we split a face, and one of the two resulting faces must be relabeled. This is a costly procedure in the initial stages of the algorithm.
In summary, our main contributions are as follows:
@@
  • Methodology: A rotation system-based algorithm for triangulating a 3D point cloud by iteratively adding edges to an initial spanning tree. With an explicit outlier classification, all the other inliers are referenced in the output mesh.
  • Automatability: An important guarantee that output mesh is always manifold enables reliable automatic workflows.
  • Generality: Extending the method to allow the creation of topological handles.
  • Controllability: The topology of output mesh can be controlled by the prescribed genus number.
  • Efficiency: To make the algorithm efficient, we use a binary tree representation of the edge loops, reducing the cost of relabeling after an edge has been added from linear to logarithmic. Moreover, we add an additional ear clipping process which allows us to avoid relabeling altogether in the later stages of the algorithm.
@@ 1.1. Related Work
Volumetric methods are often simpler to implement than combinatorial, they tend to suppress noise, and they are often good at hole closing. On the other hand, since the original points are not used directly, it is difficult to quantify the remaining noise and, more generally, the precision of the output mesh. For a comprehensive survey, the reader is referred to Berger et al. (2017), and in the following we focus on the most relevant combinatorial methods since our work falls in this category.
@@
The earliest combinatorial reconstruction method appears to be the work of Jean-Daniel Boissonnat (1984). Boissonnat proposes two algorithms, the first of which incrementally grows a polyhedron from a starting edge. The second algorithm is based on removing tetrahedra deemed exterior from the Delaunay tetrahedralization of the point cloud. The mesh resulting from this so-called sculpting process is the boundary of the union of the remaining tetrahedra. Numerous other works, notably the Crust, Power Crust, and CoCone methods of Amenta et al. (Amenta et al., 1998, 2000, 2001) and Edelsbrunner’s Wrap algorithm (Edelsbrunner, 2003) also build on 3D Delaunay tetrahedralization or its dual the Voronoi diagram. We refer the reader to Cazals and Giesen (2004) for an excellent overview. Boltcheva and Levy (2017) proposed a method in which a Voronoi cell is computed for each point and restricted to a disk that is also defined at the same point. Point connectivity can then be inferred from these restricted Voronoi cells. The method requires smoothing for real-world data but it is very efficient and easy to parallelize. More recently, Wang et al. (2022) proposed a method that is likewise based on restricted Delaunay tetrahedralization but restricts to a sequence of globally defined surfaces that approach the final output. This method appears to be effective and robust to variations in sampling density, but at the cost of making a sequence of tetrahedralizations.
@@
Bernardini et al. (1999) proposed the method known as ball-pivoting algorithm (BPA). This method is reminiscent of Boissonnat’s first method as it also grows the mesh a vertex at a time. The BPA starts from a seed triangle and uses the heuristic of pivoting a ball around an edge until it touches the next point, thereby defining a new triangle. In the presence of noise, BPA has a tendency to leave isolated points which are not touched by the rolling ball. A solution proposed by Digne et al. (2011) is to use a scale space approach where the point cloud is first smoothed and then the BPA is applied to the smoothed point cloud. This yields a reconstruction with far more of the input points, and the original vertex positions are stored and used in the final mesh.
@@
Typically, learning based methods for surface reconstruction are volumetric in nature, but combinatorial methods exist. Notably, Sharp and Ovsjanikov (2020) proposed a method which uses neural networks to iteratively propose new triangles and decide whether to include them in the output mesh. Another recent strategy employed by Sulzer et al. (2021) is to start from a Delaunay tetrahedralization and use a learning-based method to classify tetrahedra as being inside or outside. For a recent, quite broad survey of deep-learning-based surface reconstruction, the reader is referred to (Farshian et al., 2023).
@@
Not all combinatorial methods strive to include most points. A method which was expressly designed not to preserve all points was recently proposed by Zhao et al. (2023). In this work quadric error metrics are used to cluster points such that cluster representatives lie at corners and on sharp features. The mesh is then the quotient graph with respect to this clustering.
@@
The approach that is most similar to ours is the work by Robert Mencl (1995) which was later extend by Mencl and Müller (1998). Their starting point is also a minimum spanning tree, but in their case the Euclidean MST is employed whereas we use the MST of a k-nearest neighbor graph. Mencl and Müller also connect vertices and ultimately triangulate the graph, but they do not store the faces of the graph and cannot check if vertices that are connected belong to the same face. This makes it unclear whether it is feasible to avoid topological errors if the points are noisy. Unfortunately, this issue is not investigated. In later work, the same authors propose an improved method based on the so-called $\beta$ environment graphs rather than minimum spanning tree (Mencl and Müller, 2004). While the improved method is clearly effective, we are concerned that spurious handles can be introduced, especially in the presence of noise.
@@
John Robert Edmonds (1960) showed that for any graph imbued with a cyclic ordering of the incident edges for each vertex, i.e. a rotation system, there is a corresponding, topologically unique polyhedron. In the context of interactive modeling, Akleman and Chen (1999) used rotation systems to design the principles of their doubly linked face list (DLFL) representation for polyhedral shapes. Later, Akleman et al. (2003) presented a DLFL-based system where meshes are built using a minimal set of operators which includes an operator that creates a handle by adding an edge connecting two distinct faces.
2. Mesh Representation
@@
We define a graph, $G$, as the tuple $G=(V,E)$, where $V$ is a set of vertices and $E$ is a set of edges. An edge, $e=\{v,w\}$, where $v,w\in V$ is an unordered pair of vertices. For our purposes, loops where $v$ and $w$ are the same vertex will not be needed, hence $v\neq w$.
@@
The set of halfedges, $H=\{(v,w)\;|\;v,w\in V\wedge\{v,w\}\in E\}$. If there is an edge between vertices $v$ and $w$, then $(v,w)$ and $(w,v)$ both belong to $H$. Thus, halfedges, which are sometimes referred to as darts, are simply the directed versions of edges. If $v$ is the first vertex in a halfedge $(v,w)$, we say that $(v,w)$ is outgoing from $v$.
@@
A polygonal mesh, $M=(V,E,F)$ is an orientable topological polyhedron which we can obtain from a graph by associating a set of faces, $F$, with the graph. As described by Edmonds (Edmonds Jr, 1960), we can obtain the faces by associating a rotation system with the graph. A rotation system assigns cyclic orderings of the halfedges outgoing from $v$ for all $v\in V$. We can define the rotation system in terms of a function $\rho:H\rightarrow H$ which maps an outgoing halfedge, $h=(v,w)$ to the next outgoing halfedge in the cyclic ordering around $v$ as shown in Figure 2. In some cases, a vertex, $v$ has only a single incident edge and then $\rho((v,w))=(v,w)$ where $w$ is the only neighbor of $v$.
@@
In order to define the faces of a mesh in terms of a rotation system, we need the function $\iota:H\rightarrow H$ which inverts the orientation of the halfedge. In other words, if $h=(v,w)$ then $\iota(h)=(w,v)$. This means that $\iota$ is a fixed-point free involution, i.e. $\iota(\iota(h))=h$ and there is no halfedge $h$ left invariant by $\iota$.
@@
A face is defined in terms of the composition $\tau=\rho\circ\iota$. Specifically, the face associated with a halfedge, $h$, is $f_{h}=\{\tau^{n}(h)|n\in\mathbb{N}_{0}\}$ or, in other words, the orbit of $h$ under $\tau$. The set of all faces can now be defined as $F=H/\sim$ where $\sim$ is the equivalence relation defined by $h\sim h^{\prime}$ if and only if $h^{\prime}\in f_{h}$. A corner of a face, $f$, at a vertex, $v$, is given by a pair of outgoing halfedges $h$ and $h^{\prime}$ which are adjacent in the cyclic ordering, i.e. $\rho(h)=h^{\prime}$ and such that $h^{\prime}\in f$. A corner can be split by inserting a new outgoing halfedge between the two adjacent halfedges in the cyclic ordering.
@@
In the same vein as faces can be defined as orbits of $\tau$, we can see edges as the orbits of $\iota$ and vertices (including their one-rings) as the orbits of $\rho$. The actions of $\rho$, $\iota$, and $\tau$ along with their orbits are illustrated in Figure 2.
@@
Figure 2. This figure shows a partial graph consisting of vertices, edges, and halfedges (ordered pairs of vertices) indicated as fat arrows. The functions $\rho$, $\iota$, and $\tau=\rho\circ\iota$ map halfedges to halfedges. The thin arrows show the actions of the functions. Orbits of the functions are shown in red (faces), blue (edges), and green (one ring).

### v2

@@ Abstract.
Inspired by the seminal result that a graph and an associated rotation system uniquely determine the topology of a closed manifold, we propose a combinatorial method for reconstruction of surfaces from points. Our method constructs a spanning tree and a rotation system. Since the tree is trivially a planar graph, its rotation system determines a genus zero surface with a single face which we proceed to incrementally refine by inserting edges to split faces. In order to raise the genus, special handles are added in a later stage by inserting edges between different faces and thus merging them. We apply our method to a wide range of input point clouds in order to investigate its effectiveness, and we compare our method to several other surface reconstruction methods. It turns out that our approach has two specific benefits over these other methods. First, the output mesh preserves the most information from the input point cloud. Second, our method provides control over the topology of the reconstructed surface. Code is available on https://github.com/cuirq3/RsR.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/3687956journal: TOGjournalvolume: 43journalnumber: 6article: 190publicationmonth: 12submissionid: papers_693s1ccs: Computing methodologies Mesh models
@@ 1. Introduction
The ill-posed nature of the problem is due to the fact that there is generally no unique solution. Even if we restrict ourselves to only consider methods which connect all points to form a 2-manifold orientable triangle mesh, the number of possible solutions is immense for realistic input sizes. While it is trivial to connect a point to its $k$ nearest neighbors to form a graph, it is far from trivial to form a triangle mesh from this collection of edges, precisely due to the aforementioned combinatorial complexity. However, if we require the output triangle mesh to have disk topology, graph theory provides a solution that to the best of our knowledge has not been considered before. The restriction to disk topology would be problematic, but we show how it is possible to relax this restriction in order to obtain meshes with boundary curves and of arbitrary genus.
@@
Our work is based on the simple observation that a tree (in the graph theoretical sense) is always a planar graph, and the planar embedding of a tree is given by a rotation system which defines the clockwise ordering of edges incident on each vertex. Thus, if we have a spanning tree connecting all of our points and an associated rotation system, we effectively have a polygonization with a single polygon whose edges are simply the edges of the tree (with each edge of the tree appearing twice). If we imagine that the tree is actually embedded in the 2D plane, the interior of the polygon is simply the entire 2D plane minus the tree.
To obtain a triangle mesh, we recursively insert edges that split polygons into smaller polygons. The process stops when all polygons - except the boundary polygon - are triangles. A simple 2D example is shown in Figure 2 for a synthetic point cloud.
@@
For moderate input sizes and a 2D point cloud, the algorithm outlined above is simple to implement because the points are already embedded in the plane, and the rotation system is given implicitly. On the other hand, when designing an efficient and effective 3D version of the algorithm, we face several challenges. First of all, the rotation system, i.e. the ordering of outgoing edges for each vertex, must be computed consistently for all vertices. We refer the readers to Section 3.1 for the definition of a consistent rotation system. Moreover, when connecting points, several geometric considerations are important, and we want to allow connections that violate planarity in order to obtain meshes of arbitrary genus. Finally, efficiency is important: when an edge is added, we split a face, and one of the two resulting faces must be relabeled. This is a costly procedure in the initial stages of the algorithm, and the efficiency of adding edges is an important concern.
In summary, our main contributions are as follows. We
@@
  • provide a rotation system-based algorithm for triangulating a 3D point cloud by iteratively adding edges to an initial spanning tree.
  • introduce the topology test. This test checks whether the mesh would remain planar if a specific edge were added. The topology test is implemented efficiently, and it ensures both that spurious topological handles are not inserted and that non-manifold configurations are avoided.
  • allow for both topology adaptivity and control. The topology test provides us with a way to detect handle edges which we can add in order to increase the genus of the object. Since handle insertion is optional, we can also restrict the genus.
@@ 1.1. Related Work
Volumetric methods are generally simpler to implement than combinatorial methods, are usually extremely robust, and excel at hole closing. For a comprehensive survey of volumetric methods for surface reconstruction, the reader is referred to Berger et al. (2017)
Our main motivation for eschewing volumetric reconstruction in favor of combinatorial reconstruction is that the former class of methods inherently smooth the reconstructed surfaces. Admittedly, this is usually desirable, but in some cases this inherent smoothing leads to a loss of fine detail as has been noted previously (Digne et al., 2011). By decoupling the smoothing from the reconstruction, combinatorial methods leave the smoothing to downstream processing where anisotropic methods based on e.g. fuzzy median or bilateral normal smoothing (Shen and Barner, 2004; Zheng et al., 2011a) or $L_{0}$ minimization (He and Schaefer, 2013), which are designed to preserve features such as sharp edges and corners, can be brought to bear. Hence, the decoupling of smoothing and reconstruction has a tangible benefit in some cases. This benefit is only meaningful, however, if the combinatorial methods are robust. Hence, robustness is one of our key concerns.
1.1.1. Combinatorial reconstruction
@@
The earliest combinatorial reconstruction method appears to be the work of Jean-Daniel Boissonnat (1984). Boissonnat proposes two algorithms, the first of which incrementally grows a polyhedron from a starting edge. The second algorithm is based on removing tetrahedra deemed exterior from the Delaunay tetrahedralization of the point cloud. The mesh resulting from this so-called sculpting process is the boundary of the union of the remaining tetrahedra. Numerous other works, notably the Crust, Power Crust, and CoCone methods of Amenta et al. (Amenta et al., 1998, 2000, 2001) and Edelsbrunner’s Wrap algorithm (Edelsbrunner, 2003) also build on 3D Delaunay tetrahedralization or its dual the Voronoi diagram. While methods based on 3D Delaunay tetrahedralization can provide guarantees based on sampling conditions, they are generally not robust to noise. We surmise that this is because noise in the input points invariably results in a halo of tiny tetrahedra between the interior and exterior. Extracting a reasonable surface requires selecting which triangular faces to include in the output surface. Dey and Goswami (2006) provide one solution to this problem which is to select the subset of the points that lie on the boundaries of exterior Delaunay balls. However, since our aim is to include the vast majority of points, even in the presence of noise, 3D Delaunay tetrahedralization would be problematic.
Boltcheva and Levy (2017) proposed a method (Co3Ne) in which a Voronoi cell is computed for each point and restricted to a disk that is also defined at the same point. Point connectivity can then be inferred from these restricted Voronoi cells. The method requires smoothing for real-world data but it is very efficient and easy to parallelize. We compare against Co3Ne and find that it is much faster than our method but retains fewer input points and produces more holes in the output. Moreover, Co3Ne does not allow for restriction of the genus of the output shape.
More recently, Wang et al. (2022) proposed a method that is likewise based on restricted Delaunay tetrahedralization but restricts to a sequence of globally defined surfaces that approach the final output. This method appears to be effective and robust to variations in sampling density, but at the cost of making a sequence of tetrahedralizations.
We refer the reader to Cazals and Giesen (2004) for an excellent overview of (older) reconstruction methods based on Delaunay triangulation and Voronoi cells.
@@
Bernardini et al. (1999) proposed the method known as ball-pivoting algorithm (BPA). This method is reminiscent of Boissonnat’s first method as it also grows the mesh a vertex at a time. The BPA starts from a seed triangle and uses the heuristic of pivoting a ball around an edge until it touches the next point, thereby defining a new triangle. In the presence of noise, BPA has a tendency to leave isolated points which are not touched by the rolling ball. A solution proposed by Digne et al. (2011) is to use a scale space approach where the point cloud is first smoothed and then the BPA is applied to the smoothed point cloud. This yields a reconstruction with far more of the input points, and the original vertex positions are stored and used in the final mesh. Our approach includes a smoothing step similar to that introduced by Digne et al., but it is otherwise quite different. Importantly, we include even more of the input points as demonstrated by our comparisons.
@@
Typically, learning-based methods for surface reconstruction are volumetric in nature, but combinatorial methods exist. Notably, Sharp and Ovsjanikov (2020) proposed a method which uses neural networks to iteratively propose new triangles and decide whether to include them in the output mesh. Another recent strategy employed by Sulzer et al. (2021) is to start from a Delaunay tetrahedralization and use a learning-based method to classify tetrahedra as being inside or outside. Focusing on the connectivity of points, Minghua et al. (2020) applied intrinsic/extrinsic metrics to filter out undesirable connections. More recently, Rakotosaona et al. (2021) proposed a method called DSE meshing which learns selection of Delauany triangles from the input points. We compare against DSE meshing and find that it tends to be impossible to obtain a consistent orientation of the triangles. For a recent, quite broad, survey of deep-learning-based surface reconstruction, the reader is referred to (Farshian et al., 2023).
@@
By design, our method includes almost all input points in the reconstruction, but other researchers have created algorithms which intentionally reduce the number of points. A good example is the recently proposed method by Zhao et al. (2023). In this work, quadric error metrics are used to cluster points such that cluster representatives lie at corners and on sharp features. The output mesh is the quotient graph with respect to this clustering.
@@
The approach that is most similar to ours is the work by Robert Mencl (1995) which was later extended by Mencl and Müller (1998). Their starting point is also a minimum spanning tree, but in their case the Euclidean MST is employed whereas we use the MST of a k-nearest neighbor graph. Mencl and Müller also connect vertices and ultimately triangulate the graph, but they do not store the faces of the graph and, therefore, could not implement a test similar to our topology test. This means that spurious topological handles could be created since it is impossible to tell whether an inserted edge splits a face or connects different faces. For relatively noise-free point clouds this may not be much of an issue, and, unfortunately, noisy point clouds are not considered. In later work, the same authors propose an improved method based on the so-called $\beta$ environment graphs rather than minimum spanning tree (Mencl and Müller, 2004). While the improved method is clearly effective for clean point clouds, we are again concerned that topological errors could result from noise.
1.1.2. Normal estimation
Like many surface reconstruction methods, our technique requires that consistently oriented normals are associated with the input points. Fortunately, this is in itself an extensively researched area. Most of the early normal estimation methods apply plane-fitting techniques, e.g. Principal Component Analysis (PCA) (Hoppe et al., 1992; Mitra and Nguyen, 2003; Pauly et al., 2002; Huang et al., 2009), more complex surface fitting methods such as Moving Least Squares (MLS) (Alexa et al., 2001; Levin, 1998), or jet fitting (Cazals and Pouget, 2005). Recently, learning-based methods have been applied to normal estimation by directly regressing normal vectors (Guerrero et al., 2018) or fitting surfaces (Ben-Shabat and Gould, 2020; Zhu et al., 2021; Li et al., 2022). For our purposes, the older methods (e.g. (Hoppe et al., 1992) are sufficient, and in most cases we just use the normals supplied with the test data. However, we also include a test to elucidate the effect of normal noise.
1.1.3. Rotation systems
@@
John Robert Edmonds (1960) showed that for any graph imbued with a cyclic ordering of the incident edges for each vertex, i.e. a rotation system, there is a corresponding, topologically unique polyhedron [^1] Edmonds work was independent of, but much later than, the work of Lothar Heffter (1891) on which Gerhard Ringel (1965) built.. In the context of interactive modeling, Akleman and Chen (1999) used rotation systems to design the principles of their doubly linked face list (DLFL) representation for polyhedral shapes. Later, Akleman et al. (2003) presented a DLFL-based system where meshes are built using a minimal set of operators which includes an operator that creates a handle by adding an edge connecting two distinct faces.
2. Definitions
In this section, we provide a number of formal definitions which are necessary for the exposition of the method.
@@ 2. Definitions
An initial step of our approach is to form a spatially embedded graph from the 3D input points. We define a graph, $G$, as the tuple $G=(V,E)$, where $V$ is a set of vertices and $E$ is a set of edges. In the following, we will assume that a point, $\mathbf{p}_{v}\in\mathbb{R}^{3}$, is associated with each vertex $v\in V$. An edge, $e=\{v,w\}$ where $v,w\in V$, is an unordered pair of vertices. For our purposes, loops where $v$ and $w$ are the same vertex will not be needed, hence $v\neq w$.
We define a genus-0 graph as a graph that can be embedded in a manifold of genus 0 or higher such that edges intersect only at vertices. A genus-0 graph is sometimes called a planar graph since it can be drawn in the plane, but we use the term genus-0 to emphasize that an embedding surface that interpolates the vertices may be far from planar. Our method relies on the fact that if a graph is a tree, it is necessarily a genus-0 graph. This follows from the observation that starting from a single node, we can draw any tree in 2D one branch at a time such that every new branch is drawn without crossing a previously drawn branch.
A plane projectable graph (or pp graph) is a graph for which there is a 3D unit vector, $\mathbf{d}\in S^{2}$, such that when the points associated with the vertices of the graph are projected into 2D along $\mathbf{d}$, no two edges intersect except at the projected points. In this paper, we will only require that small sub-graphs of a given graph are pp graphs.
2.1. Rotation Systems
@@
In order to define rotation systems, we first need the set of halfedges, $H=\{(v,w)\;|\;v,w\in V\wedge\{v,w\}\in E\}$. If there is an edge between vertices $v$ and $w$, then $(v,w)$ and $(w,v)$ both belong to $H$. Thus, halfedges, which are sometimes referred to as darts, are simply the directed versions of edges. If $v$ is the first vertex in a halfedge $(v,w)$, we say that $(v,w)$ is outgoing from $v$.
@@
We will define a polygonal mesh as a surface constructed from a graph by associating a set of faces, $F$, with the graph. As described by Edmonds (Edmonds Jr, 1960), we can obtain the faces of a closed, two-sided polygonal mesh by associating a rotation system with the graph. A rotation system assigns cyclic orderings of the halfedges outgoing from $v$ for all $v\in V$. We can define the rotation system in terms of a bijective function $\rho:H\rightarrow H$ which maps an outgoing halfedge, $h=(v,w)$ to the next outgoing halfedge in the cyclic ordering around $v$ as shown in Figure 3. In some cases, a vertex, $v$ has only a single incident edge, and then $\rho((v,w))=(v,w)$ where $w$ is the only neighbor of $v$.
@@
In order to define the faces of a mesh in terms of a rotation system, we need the bijection $\iota:H\rightarrow H$ which inverts the orientation of the halfedge. In other words, if $h=(v,w)$ then $\iota(h)=(w,v)$. This means that $\iota$ is a fixed-point free involution, i.e. $\iota(\iota(h))=h$ and there is no halfedge $h$ left invariant by $\iota$.
@@
A face is defined in terms of the composition $\tau=\rho\circ\iota$. Specifically, $\tau$ maps a halfedge to the next halfedge in the cycle that bounds a face, and the face associated with a halfedge, $h$, is $f_{h}=\{\tau^{n}(h)|n\in\mathbb{N}_{0}\}$ or, in other words, the orbit of $h$ under $\tau$. The set of all faces can now be defined as $F=H/\sim$ where $\sim$ is the equivalence relation defined by $h\sim h^{\prime}$ if and only if $h^{\prime}\in f_{h}$. We note that in our method we define faces directly in terms of the rotation system. Only when the triangles are ouput do we convert the mesh to an indexed face set.
A corner of a face, $f$, at a vertex, $v$, is given by a pair of outgoing halfedges $h$ and $h^{\prime}$ which are adjacent in the cyclic ordering, i.e. $\rho(h)=h^{\prime}$ and such that $h^{\prime}{,\iota(h)}\in f$. A corner can be split by inserting a new outgoing halfedge between the two adjacent halfedges in the cyclic ordering.
@@
In the same vein as faces can be defined as orbits of $\tau$, we can see edges as the orbits of $\iota$ and vertices as the orbits of $\rho$. The actions of $\rho$, $\iota$, and $\tau$ along with their orbits are illustrated in Figure 3.
@@
Figure 3. This figure shows a partial graph consisting of vertices, edges, and halfedges (ordered pairs of vertices) indicated as fat arrows. The functions $\rho$, $\iota$, and $\tau=\rho\circ\iota$ map halfedges to halfedges. The thin arrows show the actions of the functions. Orbits of the functions are shown in red (faces), green (edges), and blue (one ring).